[Brahms-l] Shift report 20050617 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 17 2005 - 08:18:47 EDT
Shifter leader was AG
Shifters were 


1. There is a beam since 3:15am.<br>

2. Runs # 15560,15565-15572. Pedestal run# 15563. Non-pedestal run 15564, . Polarization measurements where done during run 15561. Run #15562 should be discarded.<br>

3. T3 channel A3-ux has demand voltage 0, voltage output 22.<br>

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 17 08:19:00 2005

This archive was generated by hypermail 2.1.8 : Fri Jun 17 2005 - 08:19:13 EDT